(a) Place the flags one standard deviation on either sid of the mean. What is the area between these two values What does the 68-95-99.7 rule say this area is? (b) Repeat for locations two and three standard deviations on either side of the mean. Again compare t 68-95-99.7 rule with the area given by the applet. 1.118 Find some proportions. Using either Table A or your calculator or software, find the proportion of observations from a standard Normal distribution that satisfies each of the following statements. In each case sketch a standard Normal curve and shade the area under the curve that is the answer to the question. (a) Z>1.75 (b) Z<1.75 (c) Z>-0.80 (d)-0.80 < Z < 1.75
Answer #1

(a)

Using z table,

P(z > 1.75) = 1-P(z < 1.75) = 1-0.9599 = 0.0401

Following is the curve:

3 2 2 3

(b)

Using z table,

P(z < 1.75) = 0.9599

Following is the curve:

3 2 2 3

(c)

Using z table,

P(z <-0.80) = 0.2119

Following is the curve:

3 2 2 3

(d)

Using z table,

P(-0.80<z<1.75)=P(zleq 1.75)-P(zleq -0.80)=0.9599-0.2119=0.7480

Following is the curve:

3 2 0 2 3